OpenStudy (anonymous):

What are the excluded values? -7 ---------- x^2-2x-15 0 and 15 3 and -5 -3 and 5

OpenStudy (asnaseer):

the excluded values are the values of x that would make the denominator zero in the fraction:\[\frac{-7}{x^2-2x-15}\]so you need to factorise the denominator first and then see which values of x would result in zero for that expression.

OpenStudy (anonymous):

so the 1st one

OpenStudy (asnaseer):

first factorise the denominator - what do you get for this?

OpenStudy (anonymous):

how do i do this

OpenStudy (asnaseer):

you might find this site helpful if you want to learn how to factorise quadratics: http://www.purplemath.com/modules/factquad.htm

OpenStudy (anonymous):

so its like (X+2)(x+15)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

then its squared

OpenStudy (asnaseer):

your factorisation is not correct as:\[(x+2)(x+15)=x(x+15)+2(x+15)=x^2+15x+2x+30=x^2+17x+30\]which is not the same as \(x^2-2x-15\).

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i need to know how to do this for a big test im taking

OpenStudy (pfenn1):

So you are looking for the factors of the expression\[x^2-2x-15\] So what number m and n exist that when we multiply\[(x+m)(x+n)=x^2-2x-15\]

OpenStudy (pfenn1):

In order for this to work, we know that m+n =-2 and mn=-15. So what two factors -15 can be added to each other to get -2?

OpenStudy (anonymous):

which one is m and which one is n

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

then

OpenStudy (pfenn1):

I'm not sure what you mean by "which one" and the "last one" But let's take a guess. If you look at the page that @asnaseer mentioned, some of this will be easier to understand.

OpenStudy (anonymous):

im saying i got the last answer

OpenStudy (anonymous):

i still dont get with that website

OpenStudy (pfenn1):

We can guess what m and n to use. Let's say m = -3 and n = 5. Then mn = -15 but m + n = 2. So that isn't right. Let's try m = 3 and n= -5. Now, mn = -15 and m+n = -2. That seems to work. So it looks like\[(x+3)(x-5)=x^2-2x-15\]. We know that the original fraction \[\frac{-7}{x^2-2x-15}\]is undefined if \[x^2-2x-15=0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont mean to rush but i need to get this question done so i can go to bed

OpenStudy (pfenn1):

(x+3)(x-5)=0 x=-3, x=5

OpenStudy (anonymous):

ok thanks

OpenStudy (pfenn1):

you're welcome